# Text Animation Illustration

### v2

@@ 4.1. Base Field and Motion Field
Learning the per-frame displacement that directly converts the input letter into animation frames is challenging. The video prior derived from foundational text-to-video models using Score Distillation Sampling (SDS) is insufficiently robust to guide the optimization, leading to severe artifacts that degrade the quality of the animation, including distortion, flickering, and abrupt appearance changes in the adjacent frame. Inspired by the CoDeF (Ouyang et al., 2023), we propose modeling the generated video in two neural displacement fields: the base field and the motion field, to address the complexity of this deformation. Both fields are represented by coordinate-based Multilayer Perceptron (MLP). To better capture high-frequency variation and represent geometry information, we project the coordinates into a higher-dimensional space using positional encoding, which is the same as the one used in NeRF (Mildenhall et al., 2021):
(3) $\small{\gamma(m)=\left(\sin(2^{0}\pi m),\cos(2^{0}\pi m),\ldots,\sin(2^{L-1}\pi m),\cos(2^{L-1}\pi m)\right)}$
$\gamma(\cdot)$ is applied separately to each dimension of the control point coordinates.
@@
The motion field, denoted as $M$, encodes the correspondence between the control points in the base shape and those in each video frame. Inspired by dynamic NeRFs (Park et al., 2021; hypernerf) and CoDeF (Ouyang et al., 2023), we represent the video as a 3D volume space, where a control point at $t$-th frame with coordinate $(x,y)$ is represented by $(x,y,t)$. Specifically, we duplicate the shared base shape $k$ times and encode $x$, $y$, and $t$ separately using Eq. 3,writing it as $P_{B}^{\prime}:\gamma(\{(P_{B},t)\}_{t=1}^{k})$. The motion field is defined as a function $M:P_{B}^{\prime}\rightarrow P_{V}$ that maps control points from the base shape to their corresponding locations $P_{V}$ in the 3D video space.
@@
To better model motion, we represent $P_{V}$ as $P_{B}+\Delta P$, focusing on learning the per-frame displacements $\Delta P=\{\Delta P^{t}\}_{t=1}^{k}$ to be applied on the base shape. Following (Gal et al., 2023), we decompose the motion into global motion (modeled by an affine transformation matrix shared by all control points of an entire frame) and local motion (predicted for each control point separately). Consider the $i$-th control point on the base shape with coordinate $(x_{B,i},y_{B,i})$, its displacement on $t$-th frame $\Delta{p}_{i}^{t}$ is summed by its local and global displacement:
(4) $\Delta{p}_{i}^{t}=\Delta p_{i}^{t,local}+\Delta{p}_{i}^{t,global}$
@@
(5) $\Delta p_{i}^{t,global}=\left[\begin{array}[]{ccc}s_{x}&sh_{x}s_{y}&d_{x}\\ sh_{y}s_{x}&s_{y}&d_{y}\\ 0&0&1\end{array}\right]\left[\begin{array}[]{ccc}\cos\theta&\sin\theta&0\\ -\sin\theta&\cos\theta&0\\ 0&0&1\end{array}\right]\cdot\begin{bmatrix}x_{B,i}\\ y_{B,i}\\ 1\end{bmatrix}-\begin{bmatrix}x_{B,i}\\ y_{B,i}\\ 1\end{bmatrix}\text{, }$
@@
where $\Delta p_{i}^{t,local}$ and all elements in the per-frame global transformation matrix are learnable.
@@
To train the base field and motion field, we distill prior knowledge from the large-scale pretrained text-to-video model, using SDS loss of Eq. 2. At each training iteration, we use a differentiable rasterizer (Li et al., 2020), denoted as $R$, to render our predicted control points set $P_{V}$ into a rasterized video (pixel format video). We proceed by selecting a diffusion timestep $\tau$, drawing a sample from a normal distribution for noise $\epsilon\sim\mathcal{N}(0,1)$, and then adding the noise to the rasterized video. The video foundation model denoise this video, based on the user prompt describing a motion pattern closely related to the word’s semantic meaning (e.g. “A camel walks steadily across the desert.” for “M” in “CAMEL”). The SDS loss is computed in Eq. 2 and guides the learning process to generate videos aligned with the desired text prompt. We jointly optimize the base field and motion field using this SDS loss. The visualized base shape demonstrates alignment with the prompt’s semantics, as shown in Fig. 5.
To maintain a legible and consistent appearance throughout the animation, we propose legibility regularization and mesh-based structure preservation regularization, which will be described in later sections.
@@
Figure 5. Base shape of “Y” for “GYM” with prompt “A man doing exercise by lifting two dumbbells in both hands”
4.2. Legibility Regularization
@@
Thus, we propose a regularization term that enforces the letter to be legible, working alongside the SDS loss to guide the optimization process. Specifically, we leverage Learned Perceptual Image Patch Similarity (LPIPS) (Zhang et al., 2018) as a loss to regularize the perceptual distance between the rasterized images of the base shape $R(P_{B})$ and the original letter $R(P_{letter})$:
(6) $\mathcal{L}_{\text{legibility}}=\text{LPIPS}\left(R\left(P_{B}\right),R\left(P_{letter}\right)\right)$
@@
4.3. Mesh-based Structure Preservation Regularization
The optimization process alters the positions of control points, sometimes leading to complex intersections between Bézier curves, as illustrated in Fig. 6(d). The rendering of Scalable Vector Graphics (SVG) adheres to the non-zero rule or even-odd rule (Foley, 1996), which determines the fill status by drawing an imaginary line from the point to infinity and counting the number of times the line intersects the shape’s boundary. The frequent intersections between bezier curves complicate the boundary, leading to alternating black and white “holes” within the image. Furthermore, these intersections between Bézier curves vary abruptly between adjacent frames, leading to severe flickering effects that degrade animation quality, see Fig. 6.
@@ 4.3. Mesh-based Structure Preservation Regularization
Figure 6. Adjacent frames of animation for letter “E” in “JET”. A large area of alternating black and white “holes” occur within each frame, as highlighted within the red circles, causing severe flickering between the adjacent frames. (d) is the visualization of frame 1, highlighting the control points and the associated Bézier curves. The illustration reveals frequent intersections among the Bézier curves leading to the flickering artifacts.
In addition, the unconstrained degrees of freedom in motion could alter the appearance of the base shape, leading to noticeable discrepancies in appearance between adjacent frames and temporal inconsistency.
@@
To address these issues, we adopt Delaunay Triangulation (Delaunay et al., 1934; Barber and Huhdanpaa, 1995) on the base shape based on control points (see Fig. 7). By maintaining the structure of the triangular mesh, we leverage the stability of triangles to prevent frequent intersections between Bézier curves, while also preserving the relative consistency of local geometry information across adjacent frames.
Specifically, we employ the angle variation (Iluz et al., 2023) of the corresponding triangular meshes in adjacent frames as a form of regularization:
(7) $\frac{1}{k\times m}\sum_{t=1}^{k}\sum_{i=1}^{m}\|\theta(T_{i,t+1})-\theta(T_{i,t})\|_{2},$
where $m$ refers to the total number of triangular meshes in each frame, $T_{i,t}$ refers to the $i$-th triangular mesh in the $t$-th frame, and $\|\theta(T_{i,t+1})-\theta(T_{i,t})\|_{2}$ refers to the sum of the squared difference in corresponding angles between two triangles. Particularly, the $(k+1)$-th frame refers to the base shape. During training, we detach the gradients of the second frame in each pair. This allows us to regularize the previous frame’s mesh structure with the next frame as a reference, and the last frame is regularized with the base shape as a reference. Hence, the structural constraint with the base shape is propagated to every frame, allowing the preservation of the geometric structure throughout the animation.
We find that this angle-based naive approach is capable of effectively maintaining the triangular structure, thereby alleviating the frequent intersections of the Beźier curves and preserving a relatively stable appearance across different frames without significantly affecting the liveliness of the motion. Furthermore, to ensure that the base shape itself mitigates the frequent intersections of Beźier curves and coarse spikes, we apply the same triangulation-based constraints between the base shape and the input letter. The whole regularization is illustrated in Fig. 7.
(8) $\begin{split}\mathcal{L}_{\text{structure}}=&\frac{1}{m}\sum_{i=1}^{m}\|\theta(T_{i,B})-\theta(T_{i,\text{letter}})\|_{2}\\ &+\frac{1}{k\times m}\sum_{t=1}^{k}\sum_{i=1}^{m}\|\theta(T_{i,t+1})-\theta(T_{i,t})\|_{2}\end{split}$
@@
Figure 7. Illustration of the Mesh-based structure preservation. We first apply this regularization between the base shape and the input letter. We propagate the structural constraint to every frame by regularizing the last frame with the base shape and regularizing every frame with its next frame.
@@
Figure 8. Visual comparisons between the baselines and our model. Text-to-image model (Gen-2) generates colorful images but fails to maintain the shape of the original letter. The pixel-based image-to-video model (DynamiCrafter) produces results with little, sometimes unreasonable motion. The general vector animation model (LiveSketch) struggles to preserve legibility and maintain a stable appearance across frames.
4.4. Frequency-based Encoding and Annealing
NeRF (Mildenhall et al., 2021) have highlighted that a heuristic application of sinusoidal functions to input coordinates, known as “positional encoding”, enables the coordinate-based MLPs to capture higher frequency content, as denoted by Eq. 3. We found that this property also applies to our MLPs that use coordinates of control points in Scalable Vector Graphics (SVG) as input. This allows the MLPs in the base and motion field to more effectively represent high-frequency information, corresponding to the detailed geometric features. Additionally, when using coordinate-based MLPs to model motion, a significant challenge is how to capture both minute and large motions. Following Nerfies (Park et al., 2021), we employ a coarse-to-fine strategy that initially targets low-frequency (large-scale) motion and progressively refines the high-frequency (localized) motions. Specifically, we use the following formula to apply weights to each frequency band $j$ in the positional encoding of the MLPs within the motion field.
(9) $w_{j}(\alpha)=\frac{1-\cos(\pi\cdot\text{clamp}(\alpha-j,0,1))}{2},$
In our experiment, this annealed frequency-based encoding resulted in higher-quality motion and detailed geometric information.
5. Experiments
@@
To comprehensively evaluate our method’s ability, we created a dataset that covers animations for all letters in the alphabet, featuring a variety of elements such as animals, humans, and objects. This dataset contains a total of 33 Dynamic Typography samples. Each sample includes a word, a specific letter within the word to be animated, and a concise text prompt describing the desired animation. We used KaushanScript-Regular as the default font.
We use text-to-video-ms-1.7b model in ModelScope (Wang et al., 2023; Luo et al., 2023) for the diffusion backbone. We apply augmentations including random crop and random perspective to all frames of the rendered videos. Each optimization takes 1000 epochs, about 40 minutes on a single H800 GPU.
@@
To illustrate our method’s capabilities, we present some generated results in Fig. 1. These animations vividly bring the specified letter to life while adhering to the prompt and maintaining the word’s readability. For further exploration, we strongly suggest the readers go through the additional examples and full-length videos on our project page.
5.1. Comparisons
@@
Within the pixel-based scenario, we compare our model against the leading text-to-video generation models Gen-2 (contributors, 2023a) (ranked first in the EvalCrafter (Liu et al., 2023) benchmark) – a commercial web-based tool, and DynamiCrafter (Xing et al., 2023), the state-of-the-art model for image-to-video generation conditioned on text. For text-to-video generation, we append the prompt with “which looks like a letter $\beta$,” where $\beta$ represents the specific letter to be animated. In the image-to-video case, we use the stylized letter generated by the word-as-image (Iluz et al., 2023) as the conditioning image.
Within the vector-based scenario, we utilize LiveSketch (Gal et al., 2023) as a framework to animate vector images. To ensure a fair comparison, we condition the animation on the stylized letter generated by the word-as-image (Iluz et al., 2023) as well.
Qualitative Comparison We present the visual comparison with baseline methods in Fig. 8. While achieving high resolution and realism, Gen-2 struggles to generate frames that keep the letter’s shape, which greatly harms the legibility. With DynamiCrafter, the “SWAN” animation exhibits minimal movement, while the “GYM” animation features unrealistic motion that deviates from the user’s prompt. Although LiveSketch can depict the user’s prompt through animation, it sacrifices legibility. Also, the letter’s appearance deteriorates throughout the animation, as demonstrated in the “SWAN” example. Our model strikes a balance between prompt-video alignment and letter legibility. It consistently generates animations that adhere to the user’s prompt while preserving the original letter’s form. This allows the animation to seamlessly integrate within the original word, as showcased by the in-context results in Fig. 8.
@@
DynamiCrafter ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{0.5151}}$ $17.8124$
@@
Ours ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathbf{0.5301}}$ ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{21.4391}}$
@@
Table 1. Quantitative results between the baselines and our model. The best and second-best scores for each metric are highlighted in red and blue respectively.
@@
Quantitative Comparison Tab. 1 presents the quantitative evaluation results. We employed two metrics, Perceptual Input Conformity (PIC) and Text-to-Video Alignment. Following DynamiCrafter (Xing et al., 2023), we computed Perceptual Input Conformity (PIC) using DreamSim’s (Poole et al., 2023) perceptual distance metric between each output frame and the input letter, averaged across all frames. This metric assesses how well the animation preserves the original letter’s appearance. To evaluate the alignment between the generated videos and their corresponding prompts (“text-to-video alignment”), we leverage the X-CLIP score (Ma et al., 2022), which extends CLIP (Radford et al., 2021) to video recognition, to obtain frame-wise image embeddings and text embeddings. The average cosine similarity between these embeddings reflects how well the generated videos align with the corresponding prompts.
@@
While Gen-2 achieves the highest text-to-video alignment score, it severely suffers in legibility preservation (lowest PIC score). Conversely, our model excels in PIC (highest score), indicating the effectiveness in maintaining the original letter’s form. While achieving the second-best text-to-video alignment score, our method strikes a balance between faithfully representing both the animation concept and the letter itself.
5.2. Ablation Study

### v3

@@ 4.1 Base Field and Motion Field
Learning the per-frame displacement that directly converts the input letter into animation frames is challenging. Directly optimizing the displacement may lead to severe artifacts that degrade the quality of the animation, including distortion, flickering, and abrupt appearance changes in the adjacent frame. Inspired by the dynamic NeRFs [43, 40] and CoDeF [38], we propose modeling the generated video in two neural displacement fields: the base field and the motion field, to address the complexity of this deformation. Both fields are represented by coordinate-based Multilayer Perceptron (MLP). To better capture high-frequency variation and represent geometry information, we project the coordinates into a higher-dimensional space using the NeRF [33] positional encoding:
$\gamma(p)=\left(\sin(2^{0}\pi p),\cos(2^{0}\pi p),\ldots,\sin(2^{L-1}\pi p),\cos(2^{L-1}\pi p)\right)\text{, }$ (3)
where $p$ refers to control point coordinates. Further details can be found in the appendix.
@@
The motion field, denoted as $M$, encodes the correspondence between the control points in the base shape and those in each video frame. Inspired by CoDeF [38], we represent the video as a 3D volume space, where a control point at $t$-th frame with coordinate $(x,y)$ is represented by $(x,y,t)$. We duplicate the shared base shape $k$ times across $k$ frames and convert the coordinates into 3D volume space, written as $P_{B}^{\prime}:\{(P_{B},t)\}_{t=1}^{k}$. The motion field is defined as a function $M:\gamma(P_{B}^{\prime})\rightarrow P_{V}$ that maps control points from the base shape to their corresponding locations $P_{V}\in\mathbb{R}^{k\times N\times 2}$ in each video frame.
@@
To better model motion, we represent $P_{V}$ as $P_{B}+\Delta P$, focusing on learning the per-frame displacements $\Delta P=\{\Delta P^{t}\}_{t=1}^{k}$ to be applied on the base shape. Following Gal et al. [14], we decompose the motion into global motion (modeled by an affine transformation matrix shared by all control points of an entire frame) and local motion (predicted for each control point separately). Consider the $i$-th control point on the base shape with coordinate $(x_{B,i},y_{B,i})$, its displacement on $t$-th frame $\Delta{p}_{i}^{t}$ is summed by its local and global displacement:
$\Delta{p}_{i}^{t}=\Delta p_{i}^{t,local}+\Delta{p}_{i}^{t,global},$ (4)
@@
$\begin{bmatrix}\Delta p_{i}^{t,global}\\ 1\end{bmatrix}=\left[\begin{array}[]{ccc}s_{x}&sh_{x}s_{y}&d_{x}\\ sh_{y}s_{x}&s_{y}&d_{y}\\ 0&0&1\end{array}\right]\left[\begin{array}[]{ccc}\cos\theta&\sin\theta&0\\ -\sin\theta&\cos\theta&0\\ 0&0&1\end{array}\right]\begin{bmatrix}x_{B,i}\\ y_{B,i}\\ 1\end{bmatrix}-\begin{bmatrix}x_{B,i}\\ y_{B,i}\\ 1\end{bmatrix},$ (5)
@@
where $\Delta p_{i}^{t,local}$ and all elements in the per-frame global transformation matrix are predicted by the MLP in the motion field.
@@
To train the base field and motion field, we distill prior knowledge from the large-scale pretrained text-to-video model using SDS computed in Eq. 2. At each training iteration, we use a differentiable rasterizer [26], denoted as $R$, to render our predicted control points set $P_{V}$ into a rasterized video (pixel format video). We select a diffusion timestep $\tau$, draw a sample from a normal distribution for noise $\epsilon\sim\mathcal{N}(0,1)$, and then add the noise to the rasterized video. The video foundation model denoise this video, based on the user prompt describing a motion pattern closely related to the word’s semantic meaning (e.g. “A camel walks steadily across the desert” for “M” in “CAMEL”). We jointly optimize the base field and motion field using the SDS loss to generate videos aligned with the desired text prompt. The visualized base shape demonstrates alignment with the prompt’s semantics, as shown in Fig. 5.
@@
Figure 5: Base shape of “Y” for “GYM” with prompt “A man doing exercise by lifting two dumbbells in both hands.”
4.2 Legibility Regularization
@@
Thus, we propose a regularization term that enforces the letter to be legible, working alongside the SDS loss to guide the optimization process. Specifically, we leverage Learned Perceptual Image Patch Similarity (LPIPS) [60] as a loss to regularize the perceptual distance between the rasterized images of the base shape $R(P_{B})$ and the original letter $R(P_{letter})$:
$\mathcal{L}_{\text{legibility}}=\text{LPIPS}\left(R\left(P_{B}\right),R\left(P_{letter}\right)\right).$ (6)
@@
4.3 Structure Preservation Regularization
The optimization process alters the positions of control points, sometimes leading to complex intersections between Bézier curves, as illustrated in Fig. 6(d). The rendering of Scalable Vector Graphics (SVG) adheres to the non-zero rule or even-odd rule [11], which determines the fill status by drawing an imaginary line from the point to infinity and counting the number of times the line intersects the shape’s boundary. The frequent intersections between Bézier curves complicate the boundary, leading to alternating black and white “holes” within the image. Furthermore, these intersections between Bézier curves vary abruptly between adjacent frames, leading to severe flickering effects that degrade animation quality, see Fig. 6. In addition, the unconstrained degrees of freedom in motion could alter the appearance of the base shape, leading to noticeable discrepancies in appearance between adjacent frames and temporal inconsistency.
@@ 4.3 Structure Preservation Regularization
Figure 6: Adjacent frames of animation for letter “E” in “JET”. A large area of alternating black and white “holes” occur within each frame, as highlighted within the red circles, causing severe flickering between the adjacent frames. (d) is the visualization of frame 1, highlighting the control points and the associated Bézier curves. The illustration reveals frequent intersections among the Bézier curves leading to the flickering artifacts.
@@
To address these issues, we adopt Delaunay Triangulation [9, 4] on the base shape based on control points, as shown in Fig. 7. By maintaining the structure of the triangular meshes, we prevent frequent intersections between Bézier curves, while preserving the relative consistency of local geometry information across adjacent frames.
@@
Figure 7: Illustration of the Mesh-based structure preservation. We first apply this regularization between the base shape and the input letter. We propagate the structural constraint to every frame by regularizing the last frame with the base shape and regularizing every frame with its next frame.
The whole regularization process is illustrated in Fig. 7. We employ the angle variation [19] of the corresponding triangular meshes in adjacent frames as regularization:
$\frac{1}{k\times m}\sum_{t=1}^{k}\sum_{i=1}^{m}\|T_{i,t+1}-T_{i,t}\|_{2}^{2},$ (7)
where $m$ refers to the total number of triangular meshes in each frame, $T_{i,t}\in\mathbb{R}^{3}$ refers to the corresponding angles in $i$-th triangular mesh in the $t$-th frame. Particularly, the $(k+1)$-th frame refers to the base shape. Consequently, the structural constraint with the base shape is propagated to every frame, allowing the preservation of the geometric structure throughout the animation. Furthermore, to ensure the base shape is geometrically similar to the input letter, we apply the same triangulation-based constraints between the base shape and the letter. $\mathcal{L}_{\text{structure}}$ can be formulated as:
$\begin{split}\mathcal{L}_{\text{structure}}=&\lambda_{1}\cdot\frac{1}{m}\sum_{i=1}^{m}\|T_{i,\text{letter}}-T_{i,B}\|_{2}^{2}\\ &+\lambda_{2}\cdot\frac{1}{k\times m}\sum_{t=1}^{k}\sum_{i=1}^{m}\|T_{i,t+1}-T_{i,t}\|_{2}^{2},\end{split}$ (8)
where $\lambda_{1}$ and $\lambda_{2}$ are weight hyperparameters .
We find that this angle-based naïve approach effectively maintains the triangular structure, thereby alleviating the frequent intersections of the Bézier curves and preserving a relatively stable appearance across different frames without significantly affecting the motion liveliness.
@@
Figure 8: Visual comparisons between the baselines and our model. Text-to-video model (Gen-2) generates colorful images but fails to maintain the shape of the original letter. The pixel-based image-to-video model (DynamiCrafter) produces results with minimal, sometimes unreasonable motion. The general vector animation model (LiveSketch) struggles to preserve legibility or maintain a consistent appearance.
5 Experiments
@@
We create a dataset via a workshop to evaluate our method’s ability. The dataset covers Dynamic Typography samples for all letters in the alphabet, featuring a variety of elements such as animals, humans, and objects, in a total of 33 samples. Each sample includes a word, a specific letter within the word to be animated, and a concise text prompt describing the desired animation. For each sample, a video with 24 frames will be generated. Each sample takes 1000 epochs for optimization, about 40 minutes on a H800 GPU.
@@
To illustrate our method’s capabilities, we present some generated results in Fig. 1. These animations vividly bring the specified letter to life while adhering to the prompt and maintaining the word’s readability. The implementation details and more results can be found in our appendix and supplementary materials.
5.1 Comparisons
@@
Within the pixel-based scenario, we compare our model against the leading text-to-video generation models Gen-2 [46] (ranked first in the EvalCrafter [27] and VBench [18] benchmark) – a commercial web-based tool, and DynamiCrafter [57], the state-of-the-art model for image-to-video generation conditioned on text. For text-to-video generation, we append the prompt with “which looks like a letter $\beta$,” where $\beta$ represents the specific letter to be animated. In the image-to-video case, we use the stylized letter generated by the Word-as-Image [19] as the conditioning image. Within the vector-based scenario, we utilize LiveSketch [14] as a framework to animate vector images. To ensure a fair comparison, we condition the animation on the stylized letter generated by the Word-as-Image [19] as well. For fairness, all animations to be compared are composed of 24 frames, each lasting 3 seconds, and are rendered at a resolution of $256\times 256$.
Qualitative Comparison We present two samples for visual comparison in Fig. 8. For more comparisons and full videos, please check our supplementary materials. While achieving high resolution and realism, Gen-2 struggles to generate frames that keep the letter’s shape, which greatly harms the legibility. With DynamiCrafter, the “SWAN” animation exhibits minimal movement, while the “GYM” animation features unrealistic motion that deviates from the user’s prompt. Although LiveSketch can depict the user’s prompt through animation, it sacrifices legibility. Also, the animated letter fails to maintain a stable appearance throughout the animation, as demonstrated in the “SWAN” example. Our model strikes a balance between prompt-video alignment and letter legibility. It consistently generates animations that adhere to the user’s prompt while preserving the original letter’s form. This allows the animation to seamlessly integrate within the original word, as showcased by the in-context results in Fig. 8.
@@
DynamiCrafter $0.5151$ $17.8124$
@@
Ours ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathbf{0.5301}}$ $21.4391$
@@
Table 1: Quantitative results between the baselines and our model. The best score for each metric is highlighted in red. Our model gets the best score in PIC, and second-best score in text-to-video alignment, indicating a balance between faithfully representing the user’s prompt and maintaining the legibility.
@@
Quantitative Comparison Tab. 1 presents the quantitative evaluation results. We employ two metrics, Perceptual Input Conformity (PIC) and Text-to-Video Alignment. Following DynamiCrafter [57], we compute Perceptual Input Conformity (PIC) using DreamSim’s [42] perceptual similarity metric between each output frame and the input letter, averaged across all frames. This metric assesses how well the animation preserves the original letter’s appearance. To evaluate the alignment between the generated videos and their corresponding prompts (Text-to-Video Alignment), we leverage the X-CLIP score [30], which extends CLIP [44] to video recognition, to obtain frame-wise image embeddings and text embeddings. The average cosine similarity between these embeddings reflects how well the generated videos align with the corresponding prompts.
@@
While Gen-2 achieves the highest text-to-video alignment score, it severely suffers in legibility preservation (lowest PIC score). Conversely, our model excels in PIC, indicating the effectiveness in maintaining the original letter’s form. Our model also achieves the second-best text-to-video alignment score, indicating it faithfully depicts the animation prompt while preserving legibility.
5.2 Ablation Study